# Promote, Suppress, Iterate: How Language Models Answer One-to-Many Factual Queries

### v2

@@ Abstract
To answer one-to-many factual queries (e.g., listing cities of a country), a language model (LM) must simultaneously recall knowledge and avoid repeating previous answers. How are these two subtasks implemented and integrated internally? Across multiple datasets and models, we identify a promote-then-suppress mechanism: the model first recalls all answers, and then suppresses previously generated ones. Specifically, LMs use both the subject and previous answer tokens to perform knowledge recall, with attention propagating subject information and MLPs promoting the answers. Then, attention attends to and suppresses previous answer tokens, while MLPs amplify the suppression signal. Our mechanism is corroborated by extensive experimental evidence: in addition to using early decoding and causal tracing, we analyze how components use different tokens by introducing both Token Lens, which decodes aggregated attention updates from specified tokens, and a knockout method that analyzes changes in MLP outputs after removing attention to specified tokens. Overall, we provide new insights into how LMs’ internal components interact with different input tokens to support complex factual recall.  [^1] Code is available at https://github.com/Lorenayannnnn/how-lms-answer-one-to-many-factual-queries.
@@ 1 Introduction
To understand the overall process, we early decode (Nostalgebraist, 2020) the output of attention and MLPs to examine how the logits of the subject and answer tokens change across layers. We find that LMs first promote all answers and then suppress the ones that have been previously generated. Specifically, attention copies the subject information at the middle layers and MLPs promote all possible answers. Then, both components suppress previous answer tokens at late layers. These observations hold for both Llama-3-8B-Instruct and Mistral-7B-Instruct across three datasets.
@@ 2 Related Work
Recent works have studied the functions of MLPs and attention in knowledge recall given subject-relation pairs. Geva et al. (2023) and Meng et al. (2022) demonstrate that MLPs enrich subject representations at early layers, while Geva et al. (2022) and Merullo et al. (2023b) highlight how MLPs promote correct answer tokens by writing updates to the residual stream and adjusting the vocabulary probabilities. This mechanism is still essential for the model to generate multiple answers tied to the given subject. Prior works have also shown that attention plays a key role in extracting important tokens and suppressing repeated ones (Wang et al., 2022; McDougall et al., 2023; Merullo et al., 2023a; Tigges et al., 2024), which is essential for preventing the model from generating duplicate answers. Merullo et al. (2024) further decomposes attention heads and identifies low-rank subspaces in which components communicate to selectively inhibit repetitive items from a list given in the context, which also involves list processing and repetition avoidance but not recalling factual knowledge from model parameters.
@@ 3.1 Task: One-to-Many Knowledge Recall
[…]
@@ 3.2 Datasets and Models
Dataset Subject ($s$) Relation ($r$) Object ($o$) # Entries Llama-3-8B-Instruct (Acc) Mistral-7B-Instruct (Acc)
Country-Cities Country contains Cities 168 122/168 (72.6%) 116/168 (69.0%)
Artist-Songs Artist performer of Songs 2077 261/2077 (12.6%) 206/2077 (9.9%)
Actor-Movies Actor acted in Movies 8790 1285/8790 (14.6%) 965/8790 (11.0%)
Table 1: Models’ performance on all three datasets. Lower accuracy may be attributed to long-tail entities, outdated data, and overly strict exact-match evaluations. Our analysis focuses on the correct cases.
We curate three 1MKR datasets on different topics: (1) cities of a country, [^2] https://simplemaps.com/data/world-cities (2) songs performed by an artist, [^3] https://www.kaggle.com/datasets/salvatorerastelli/spotify-and-youtube, and (3) movies acted in by an actor or actress. [^4] https://www.kaggle.com/datasets/darinhawley/imdb-films-by-actor-for-10k-actors A summary of the datasets is provided in Table 1. For each dataset, we filter out subjects that are associated with fewer than three object entities for the specified relation.
We study two LMs: Llama-3-8B-Instruct AI@Meta (2024) and Mistral-7B-Instruct AI (2024). Refer to Appendix A for the prompt templates we use for each model and dataset. To create the data for analyzing LMs’ behaviors, we first generate three answers using greedy decoding, ensuring consistent outputs for examining component behaviors across different answer steps. We then retain the entries where all three predicted answers are correct to focus on cases where the models’ knowledge is accurate.
Table 1 shows the number of correct predictions made by the models across the datasets. The low accuracy may be explained by (1) long-tail entities (e.g., less popular actors or songs), (2) outdated datasets compared to the model’s knowledge, and (3) the strict use of exact match evaluation (e.g., “Mission: Impossible” is considered incorrect even if given “Mission: Impossible - Fallout” is in the label list). For all (dataset, model) pairs, we have at least 100 correct instances, providing a sufficient sample size for the analysis. For the rest of the paper, we focus only on the correct cases. When analyzing models’ behaviors at step $i$ ($i=1,2,3$), we keep all tokens before the first token of the $i$th answer as input. Refer to Appendix B for examples and details. We report results macro-averaged across both models and all datasets in the main section. Refer to the appendix for full results of all answer steps and specific models and datasets. We run all experiments on a single RTX A6000 GPU.
@@ 4.2 LMs Promote Then Suppress
Figure 2: Logit of the subject and answer tokens from unembedding attention and MLP outputs. Attention primarily promotes the subject at the middle layers, then promotes new answers and suppresses previous answers at deeper layers. MLPs consistently promote all answers; at deeper layers, they also decrease the logits of previously generated answers. Early layers are omitted as logits are near zero. See Appendix C for full figures.
@@
As shown in Figure 2, attention outputs positive logits for the subject token in the middle layers across all three answer steps. While the three answers are slightly promoted at layer $25$, their logits are still close to zero and have a smaller magnitude compared to that of the subject at the middle layers. This pattern indicates that attention copies or propagates subject information at the last token position. Interestingly, the answer promotion pattern is more evident in the Country-Cities dataset (Figure 9, Figure 10) but not in the Artist-Songs and the Actor-Movies datasets (Figure 11, Figure 12, Figure 13, Figure 14).
@@
From the middle to later layers, MLPs consistently output positive logits for all three possible answers (Figure 2). These logits increase across generation steps, with their magnitude significantly exceeding that of attention logits. These findings suggest that MLPs strongly promote all possible answers regardless of prior predictions, thereby providing a stronger answer promotion signal compared to attention.
@@
Both attention and MLPs suppress answers that have been generated previously. Starting from layer $28$, attention outputs negative logits for $o^{(1)}$ at step $2$ and for both $o^{(1)}$ and $o^{(2)}$ at step $3$ (Figure 2). Similarly, MLPs decrease the logit of previous answers at the same layer. Since MLPs themselves cannot attend back to early tokens, this suppression likely results from leveraging suppression signals from attention, a hypothesis further investigated in Section 6.3.
@@ 5 Which Tokens Matter?
To better understand the promote-then-suppress mechanism, we now investigate how LMs implement knowledge recall and repetition avoidance. In this section, we use causal tracing Meng et al. (2022) to identify the input tokens that most influence model predictions. In Section 6, we analyze how these tokens are used by attention and MLPs to facilitate knowledge recall and repetition avoidance.
@@ 5.1 Which Tokens Should Be Noised?
Prior work shows that in order to recall knowledge, LMs encode information about relevant object entities in subject tokens and retrieve this information via attention (Geva et al., 2023; Meng et al., 2022). Other work shows that LMs avoid repetition by using attention heads to attend to previous tokens and suppress them (McDougall et al., 2023; Wang et al., 2022; Merullo et al., 2023a). Thus, we hypothesize that the subject and previous answer tokens play decisive roles in our two key sub-tasks (Section 3.1).
@@
Figure 3: The impact of attention and MLPs’ activations on LMs’ predictions when intervening on the subject (left) and previous answer tokens (right) at step $2$. The probability differences all peak around or above $0.55$, reflecting the importance of both the subject and previous answer tokens. See Appendix D for figures of other answer steps, which have similar patterns.
@@
Figure 3 visualizes the impact of attention and MLPs on LMs’ predictions when intervening on the subject tokens at step $2$ (Refer to Appendix D for figures of other answer steps and specific models and datasets, which have similar patterns). The probability difference peaks around or above $0.55$ for both components, confirming our hypothesis that the subject plays a crucial role in knowledge recall. Attention’s contributions peak in the middle layers at the last token, while MLPs dominate in early layers at the subject token and in late layers at the last token. These observations suggest that attention propagates subject information from early MLP layers to the last token, where MLPs may leverage it for answer promotion, as discussed in Section 6.2.
@@
Noising previous answer tokens also leads to high probability changes in LMs’ output probabilities, with an average difference of around or above $0.55$ across answer steps (Figure 3). This finding supports our hypothesis that previous answer tokens are also critical to LMs’ outputs. Similar to the results of noising the subject, attention’s contributions peak in both the middle and the last layers at the last token. MLPs dominate in early layers at the previous answer positions and in late layers at the last token, reflecting that the previous answer tokens are used by both components to make nontrivial contributions to models’ predictions.
@@ 6.1 Methodology for Analyzing Tokens
Finally, following the same approach of (early) decoding attention and MLP outputs in Section 4.1, we unembed $a^{(l)}_{e}$ to obtain the logits of the first token of the subject and answers and examine how attention uses the target tokens to perform promotion or suppression.
@@
Let $m^{(l)}$ and $m^{\prime(l)}$ denote the MLP output at layer $l$ before and after applying the attention knockout respectively. We unembed these outputs using the same early decoding approach described in Section 4.1. By subtracting the logits derived from $m^{\prime(l)}$ from those of $m^{(l)}$, we examine the difference in the logits of the subject and the answer tokens. A positive difference value indicates MLPs use the knocked-out tokens to promote a token; a negative difference means suppression.
@@ 6.2 Role of Subject Tokens
Figure 4: Token Lens logit values (left) and MLP logit differences (right) of subject and answer tokens when attending to or knocking out the subject tokens. Attention promotes and extracts subject information in the middle layers but suppresses it in later layers. MLPs promote the answers and suppress the subject at deeper layers.
@@
As shown in Figure 4, attention to the subject greatly increases the subject token’s logit at the middle layers. To a lesser degree, it also promotes answer tokens, particularly at layer $25$. [^7] Thus, the observation from Section 4.2 that attention promotes answers at layer $25$ can be attributed to the subject token. Answer promotion is most pronounced in the Country-Cities dataset (Figure 23, Figure 24) but less evident in the Artist-Songs and Actor-Movies datasets (Figure 25, Figure 26, Figure 27, Figure 28). In all datasets, the subject logit is still larger than that of each answer across all answer steps, demonstrating that attention primarily copies or propagates subject information from the subject to the last token position.
@@ 6.3 Role of Previous Answer Tokens
As shown in Figure 5, all previous answer tokens have negative MLP logit differences at late layers. For instance, $o_{1}$ has negative logits at step $2$ starting around layer $27$; $o_{1}$ and $o_{2}$ exhibit similar patterns at step $3$. This suppression aligns with attention’s role in inhibiting previously generated tokens, suggesting that MLPs amplify these suppression signals to prevent repetition.
@@
Surprisingly, we observe positive MLP logit differences for new answers across answer steps (Figure 5). Specifically, the logit differences of both $o_{2}$ and $o_{3}$ are positive when intervening on $o_{1}$ at step $2$; $o_{3}$ has positive logits differences when intervening on $o_{1}$ or $o_{2}$ at step $3$. This pattern shows that MLPs also leverage previous answer tokens to promote new answers. Since LMs already promote all relevant answers when predicting previous answers, it is plausible that the models reuse these prior computations to promote new answers. These findings show that the subject token is not the sole source of answer promotion (Section 5.1). The previous answer tokens also have a positive (but smaller) effect on answer promotion.
@@ 6.4 Role of Last Token
Starting from layer $28$, attention from the last token to itself significantly increases the logit of all three answers (Figure 6). At each step $i$, the logit for the answer ${o^{(i)}}$ is consistently the highest among the three answers. This result suggests that attention at the last token aggregates information from earlier layers related to knowledge recall and suppression, preparing the model for generating the next prediction.
@@
Interestingly, we observe MLPs output negative logit differences for the subject and all three answers in the final layers when knocking out the attention from the last token to itself (Figure 6). The answer ${o^{(i)}}$ for each step $i$ consistently has the most negative logit differences. In other words, without having access to the attention output of the last token, MLPs output even higher logits for the subject and the answers. This behavior suggests a backup mechanism: without direct attention to the last token that aggregates information from early input tokens, the model may not have sufficient promotion and differentiation of the three answers. MLPs compensate this by further promoting the three answers to encourage the predictions to be correct. Similarly, Wang et al. (2022) find backup token mover attention heads that become active when the original token mover heads are ablated.
7 Conclusion
Future Work.
Future work could investigate possible redundancies in the model, as multiple tokens—such as the subject and previous answers—contribute to promoting new answers. This result raises the question of whether LMs redundantly encode knowledge and if it is necessary. Additionally, our analyses only focus on the correct cases. Examining the patterns when LMs hallucinate could provide insights for mitigating such errors.
Limitations
Acknowledgments
Appendix A Prompt Templates
Following are the prompt templates that we use for each model and dataset. Llama-3-8B-Instruct:
  • Country-Cities: List three cities from country:
  • Artist-Songs: List three songs performed by artist:
  • Actor-Movies: List three movies acted by actor:
Mistral-7B-Instruct:
  • Country-Cities: List the name of three cities from country:
  • Artist-Songs: List three songs performed by artist:
  • Actor-Movies: List the name of three movies acted by actor:
Appendix B Sample Responses and Example of Analysis Data Creation
  • List three cities from China: 1. Beijing 2. Shanghai 3. Guangzhou
Mistral-7B-Instruct:
[…]
We filter out responses that contain incorrect object entities and only focus on the correct cases for analyses. For the artist-songs dataset, we use the Spotify API  [^8] https://developer.spotify.com/documentation/web-api to extend the song lists and keep them more up-to-date. To create data for analyzing, for example, Mistral-7B-Instruct’s behavior when predicting the first answer about Ed Sheeran, we will use "List the name of three songs performed by Ed Sheeran: 1." as the input and examine models’ behavior when predicting "Shape".
Appendix C Decoding Attention and MLP Outputs Results
Figure 7 and Figure 8 are the full figures of logit of the subject and target entity tokens from decoding attention and mlp output across layers and answer steps. Figure 9, Figure 10, Figure 11, Figure 12, Figure 13, Figure 14 are the figures for specific models and datasets. As can be seen from Figure 9 and Figure 10, attention performing answer promotion at middle layers is more evident in the Country-Cities dataset. However, it is much less evident in the other two datasets (Figure 11, Figure 12, Figure 13, Figure 14). Refer to https://drive.google.com/drive/folders/1Xnk3lPLuqjmNABfrJvcJ4mSM9EBvYoub?dmr=1&ec=wgc-drive-globalnav-goto for the figures without early layers omitted.
Figure 9: Attention and MLP output logits of Llama-3-8B-Instruct on Country-Cities dataset.
Figure 10: Attention and MLP output logits of Mistral-7B-Instruct on Country-Cities dataset.
Figure 11: Attention and MLP output logits of Llama-3-8B-Instruct on Artist-Songs dataset.
Figure 12: Attention and MLP output logits of Mistral-7B-Instruct on Artist-Songs dataset.
Figure 13: Attention and MLP output logits of Llama-3-8B-Instruct on Actor-Movies dataset.
Figure 14: Attention and MLP output logits of Mistral-7B-Instruct on Actor-Movies dataset.
Appendix D Causal Tracing Results
Figure 15 and Figure 16 are the full figures for causal tracing when noising the subject and previous answer tokens across all three answer steps. Refer to https://drive.google.com/drive/folders/1aG-GZEIZ_EgUKQ8Vhe_Lv0mHILxIZfms?dmr=1&ec=wgc-drive-globalnav-goto for figures of specific models and datasets.
Figure 15: The impact of attention and MLPs’ activations on LMs’ predictions when intervening on the subject tokens across three answer steps macro-averaged across all models and datasets. Attention contributions dominate in the middle layers at the last token, while MLPs are important in early layers at the subject token and in late layers at the last token. The probability differences all peak around or above $0.55$, reflecting the importance of the subject tokens.
Figure 16: The impact of attention and MLPs’ activations on LMs’ predictions when intervening on previous answer tokens at step $2$ and $3$ macro-averaged across all models and datasets. Attention is important in both the middle and the last layers at the last token position. MLPs’ contributions are critical in early layers at the previous answer positions and in final layers at the last token. The probability differences all peak around or above $0.54$, indicating previous answer tokens are critical to models’ predictions.
Appendix E Critical Token Analysis Results
Figure 17, Figure 18, Figure 19, Figure 20, Figure 21, Figure 22 are the complete results for Token Lens and Attention Knockout analyses on the subject token, previous answer tokens, and the last token. The results are macro-averaged across three answer steps and aggregated over all models and datasets. Figure 23, Figure 24, Figure 25, Figure 26, Figure 27, Figure 28 are the Token Lens and attention Knockout results on the subject token from different models and datasets. The pattern of attention using the subject token to promote answers is more prominent in the Country-Cities dataset (Figure 23, Figure 24) compared to the other two datasets (Figure 25, Figure 26, Figure 27, Figure 28). Refer to https://drive.google.com/drive/folders/1HtMtgm63ZZDfAnjeFDLJqMwSvLSyWAlj?dmr=1&ec=wgc-drive-globalnav-goto for dataset- and model-specific figures on all different tokens without early layers omitted.
Figure 17: Token Lens logit values of subject and answer tokens across layers and answer steps when attending to the subject (macro-averaged across all datasets and models). Attention promotes and extracts subject information in the middle layers while suppressing it in later layers.
Figure 18: Average logit differences of the subject and answer tokens between MLP outputs with and without knocking out attention from the last to the subject tokens. Positive logit differences for the answers and negative differences for the subject in later layers show that MLPs use the subject information to promote answers and suppress the subject.
Figure 19: Token Lens logit values subject and answer tokens across layers and answer steps $2$ and $3$ (macro-averaged across all datasets and models) when attending to previous answers. The logit of the attended answer is negative at later layers, showing that the attention is suppressing previously generated answers.
Figure 20: Average logit differences for subject and answer tokens between MLP outputs with and without knocking attention from the last to previous answer tokens. All previously generated answer tokens have negative logits, and all new answers have positive logits. This result suggests that MLPs use previous answers for both repetition suppression and new answer promotion.
Figure 21: Token Lens logit values of subject and answer tokens across layers and answer steps when attending to the last token (macro-averaged across all datasets and models). Attention promotes all three answers and the subject at the final layers, with the answer for the current step having the highest logit.
Figure 22: Average logit differences for subject and answer tokens between MLP outputs with and without knocking attention from the last token to itself. The logit differences of all three answers and the subject are negative at the late layers, meaning MLPs output higher logits when it does not have information from the last token. This pattern may suggest a compensation behavior for the absence of direct attention to the last token to encourage the outputs to still be correct.
Figure 23: Token Lens logit values (left) and MLP logit differences (right) of subject and answer tokens of Llama-3-8B-Instruct on Country-Cities dataset when attending to or knocking out the subject tokens.
Figure 24: Token Lens logit values (left) and MLP logit differences (right) of subject and answer tokens of Mistral-7B-Instruct on Country-Cities dataset when attending to or knocking out the subject tokens.
Figure 25: Token Lens logit values (left) and MLP logit differences (right) of subject and answer tokens of Llama-3-8B-Instruct on Artist-Songs dataset when attending to or knocking out the subject tokens.
Figure 26: Token Lens logit values (left) and MLP logit differences (right) of subject and answer tokens of Mistral-7B-Instruct on Artist-Songs dataset when attending to or knocking out the subject tokens.
Figure 27: Token Lens logit values (left) and MLP logit differences (right) of subject and answer tokens of Llama-3-8B-Instruct on Actor-Movies dataset when attending to or knocking out the subject tokens.
Figure 28: Token Lens logit values (left) and MLP logit differences (right) of subject and answer tokens of Mistral-7B-Instruct on Actor-Movies dataset when attending to or knocking out the subject tokens.
Appendix F Are Knowledge Recall and Suppression Independent?

### v3

@@ Abstract
To answer one-to-many factual queries (e.g., listing cities of a country), a language model (LM) must simultaneously recall knowledge and avoid repeating previous answers. How are these two subtasks implemented and integrated internally? Across multiple datasets, models, and prompt templates, we identify a promote-then-suppress mechanism: the model first recalls all answers, and then suppresses previously generated ones. Specifically, LMs use both the subject and previous answer tokens to perform knowledge recall, with attention propagating subject information and MLPs promoting the answers. Then, attention attends to and suppresses previous answer tokens, while MLPs amplify the suppression signal. Our mechanism is corroborated by extensive experimental evidence: in addition to using early decoding and causal tracing, we analyze how components use different tokens by introducing both Token Lens, which decodes aggregated attention updates from specified tokens, and a knockout method that analyzes changes in MLP outputs after removing attention to specified tokens. Overall, we provide new insights into how LMs’ internal components interact with different input tokens to support complex factual recall.  [^1] Code is available at https://github.com/Lorenayannnnn/how-lms-answer-one-to-many-factual-queries.
@@ 1 Introduction
To understand the overall process, we early decode (Nostalgebraist, 2020) the output of attention and MLPs to examine how the logits of the subject and answer tokens change across layers. We find that LMs first promote all answers and then suppress the ones that have been previously generated. Specifically, attention copies the subject information at the middle layers and MLPs promote all possible answers. Then, both components suppress previous answer tokens at late layers. These observations hold for both Llama-3-8B-Instruct and Mistral-7B-Instruct-v0.2 across three datasets.
@@ 2 Related Work
Recent works have studied the functions of MLPs and attention in knowledge recall given subject-relation pairs. Meng et al. (2022) and Geva et al. (2023) demonstrate that MLPs enrich subject representations at early layers, while Geva et al. (2022) and Merullo et al. (2023b) highlight how MLPs promote correct answer tokens by writing updates to the residual stream and adjusting the vocabulary probabilities. This mechanism is still essential for the model to generate multiple answers tied to the given subject. Prior works have also shown that attention and MLPs play a key role in extracting important tokens and suppressing repeated ones (Wang et al., 2022; McDougall et al., 2023; Voita et al., 2023; Merullo et al., 2023a; Tigges et al., 2024), which is essential for preventing the model from generating duplicate answers. Merullo et al. (2024) further decomposes attention heads and identifies low-rank subspaces in which components communicate to selectively inhibit repetitive items from a list given in the context, which also involves list processing and repetition avoidance but not recalling factual knowledge from model parameters.
@@ 3.1 Task: One-to-Many Knowledge Recall
In 1MKR, a language model is given a subject entity $s$ and a relation $r$, and must generate a set of corresponding object entities $O=\{o^{(1)},o^{(2)},\dotsc,o^{(n)}\}$ that are related to $s$ through $r$. All generated object entities must be distinct, that is, $o^{(i)}\neq o^{(j)}$ for $i\neq j$. For example, given $s=\text{"U.S.A."}$ and $r=\text{"cities of"}$, one possible valid set of object entities is $O=\{\text{Los Angeles},\text{San Francisco},\text{Seattle}\}$. To perform this task, the model must perform two key subtasks:
@@ 3.2 Datasets and Models
Dataset Subject ($s$) Relation ($r$) Object ($o$) # Entries Llama-3-8B-Instruct (Acc) Mistral-7B-Instruct-v0.2 (Acc)
Country-Cities Country contains Cities 168 122/168 (72.8%) 118/168 (70.2%)
Artist-Songs Artist performer of Songs 2077 276/2077 (13.3%) 221/2077 (10.6%)
Actor-Movies Actor acted in Movies 8790 1235/8790 (14.1%) 799/8790 (9.1%)
Table 1: Models’ performance on all three datasets averaged across three prompt templates. Lower accuracy may be attributed to long-tail entities, outdated data, and overly strict exact-match evaluations. Our analysis focuses on the correct cases. See Appendix˜A for the prompt templates and per-template performance.
We curate three 1MKR datasets on different topics: (1) cities of a country, [^2] https://simplemaps.com/data/world-cities (2) songs performed by an artist, [^3] https://www.kaggle.com/datasets/salvatorerastelli/spotify-and-youtube, and (3) movies acted in by an actor or actress. [^4] https://www.kaggle.com/datasets/darinhawley/imdb-films-by-actor-for-10k-actors A summary of the datasets is provided in Table˜1. For each dataset, the number of object entities n=3n=3.555We also tested larger $n$. The models fail to generate at least $100$ correct cases except for the Actor-Movies dataset with $n=5$, where all major results align with those discussed in the main section. See Appendix F for details. We filter out subjects that are associated with fewer than three object entities for the specified relation.
We study two LMs: Llama-3-8B-Instruct AI@Meta (2024) and Mistral-7B-Instruct-v0.2 AI (2024). We have three prompt templates for each model and dataset, which are shown in Appendix˜A. To create the data for analyzing LMs’ behaviors, we first generate three answers using greedy decoding, ensuring consistent outputs for examining component behaviors across different answer steps. We then retain the entries where all three predicted answers are correct to focus on cases where the models’ knowledge is accurate.
Table˜1 shows the number of correct predictions made by the models across the datasets. The low accuracy may be explained by (1) long-tail entities (e.g., less popular actors or songs), (2) outdated datasets compared to the model’s knowledge, and (3) the strict use of exact match evaluation (e.g., “Mission: Impossible” is considered incorrect even if given “Mission: Impossible - Fallout” is in the label list). For all (dataset, model) pairs, we have at least 100 correct instances, providing a sufficient sample size for the analysis. For the rest of the paper, we focus only on the correct cases. When analyzing models’ behaviors at step $i$ ($i=1,2,3$), we keep all tokens before the first token of the $i$th answer as input. Refer to Appendix˜B for examples and details. We report results macro-averaged across all models, datasets, and prompt templates in the main section. Refer to the appendix for full results of all answer steps and specific models and datasets. We run all experiments on a single RTX A6000 GPU.
@@ 4.2 LMs Promote Then Suppress
Figure 2: Logit of the subject and answer tokens from unembedding attention and MLP outputs. Attention primarily promotes the subject at the middle layers, then promotes new answers and suppresses previous answers at deeper layers. MLPs consistently promote all answers; at deeper layers, they also decrease the logits of previously generated answers. Early layers are omitted as logits are near zero. See Appendix˜C for full figures.
@@
As shown in Figure˜2, attention outputs positive logits for the subject token in the middle layers across all three answer steps. While the three answers are slightly promoted at layer $25$, their logits are still close to zero and have a smaller magnitude compared to that of the subject at the middle layers. This pattern indicates that attention copies or propagates subject information at the last token position. Interestingly, the answer promotion pattern is more evident in the Country-Cities dataset (Figure˜10, Figure˜11) but not in the Artist-Songs and the Actor-Movies datasets (Figure˜12, Figure˜13, Figure˜14, Figure˜15).
@@
From the middle to later layers, MLPs consistently output positive logits for all three possible answers (Figure˜2). These logits increase across generation steps, with their magnitude significantly exceeding that of attention logits. These findings suggest that MLPs strongly promote all possible answers regardless of prior predictions, thereby providing a stronger answer promotion signal compared to attention.
@@
Both attention and MLPs suppress answers that have been generated previously. Starting from layer $28$, attention outputs negative logits for $o^{(1)}$ at step $2$ and for both $o^{(1)}$ and $o^{(2)}$ at step $3$ (Figure˜2). Similarly, MLPs decrease the logit of previous answers at the same layer. Since MLPs themselves cannot attend back to early tokens, this suppression likely results from leveraging suppression signals from attention, a hypothesis further investigated in Section˜6.3.
@@ 5 Which Tokens Matter?
To better understand the promote-then-suppress mechanism, we now investigate how LMs implement knowledge recall and repetition avoidance. In this section, we use causal tracing Meng et al. (2022) to identify the input tokens that most influence model predictions. In Section˜6, we analyze how these tokens are used by attention and MLPs to facilitate knowledge recall and repetition avoidance.
@@ 5.1 Which Tokens Should Be Noised?
Prior work shows that in order to recall knowledge, LMs encode information about relevant object entities in subject tokens and retrieve this information via attention (Geva et al., 2023; Meng et al., 2022). Other work shows that LMs avoid repetition by using attention heads to attend to previous tokens and suppress them (McDougall et al., 2023; Wang et al., 2022; Merullo et al., 2023a). Thus, we hypothesize that the subject and previous answer tokens play decisive roles in our two key sub-tasks (Section˜3.1).
@@
Figure 3: The impact of attention and MLPs’ activations on LMs’ predictions when intervening on the subject (left) and previous answer tokens (right) at step $2$. The probability differences all peak around or above $0.55$, reflecting the importance of both the subject and previous answer tokens. See Appendix˜D for figures of other answer steps, which have similar patterns.
@@
Figure˜3 visualizes the impact of attention and MLPs on LMs’ predictions when intervening on the subject tokens at step $2$ (Refer to Appendix˜D for figures of other answer steps and specific models and datasets, which have similar patterns). The probability difference peaks around or above $0.55$ for both components, confirming our hypothesis that the subject plays a crucial role in knowledge recall. Attention’s contributions peak in the middle layers at the last token, while MLPs dominate in early layers at the subject token and in late layers at the last token. These observations suggest that attention propagates subject information from early MLP layers to the last token, where MLPs may leverage it for answer promotion, as discussed in Section˜6.2.
@@
Noising previous answer tokens also leads to high probability changes in LMs’ output probabilities, with an average difference of around or above $0.55$ across answer steps (Figure˜3). This finding supports our hypothesis that previous answer tokens are also critical to LMs’ outputs. Similar to the results of noising the subject, attention’s contributions peak in both the middle and the last layers at the last token. MLPs dominate in early layers at the previous answer positions and in late layers at the last token, reflecting that the previous answer tokens are used by both components to make nontrivial contributions to models’ predictions.
@@ 6.1 Methodology for Analyzing Tokens
Finally, following the same approach of (early) decoding attention and MLP outputs in Section˜4.1, we unembed $a^{(l)}_{e}$ to obtain the logits of the first token of the subject and answers and examine how attention uses the target tokens to perform promotion or suppression.
@@
Let $m^{(l)}$ and $m^{\prime(l)}$ denote the MLP output at layer $l$ before and after applying the attention knockout respectively. We unembed these outputs using the same early decoding approach described in Section˜4.1. By subtracting the logits derived from $m^{\prime(l)}$ from those of $m^{(l)}$, we examine the difference in the logits of the subject and the answer tokens. A positive difference value indicates MLPs use the knocked-out tokens to promote a token; a negative difference means suppression.
@@ 6.2 Role of Subject Tokens
Figure 4: Token Lens logit values (left) and MLP logit differences (right) of subject and answer tokens when attending to or knocking out the subject tokens. Attention promotes and extracts subject information in the middle layers but suppresses it in later layers. MLPs promote the answers and suppress the subject at deeper layers. Refer to Appendix˜E for full figures.
@@
As shown in Figure˜4, attention to the subject greatly increases the subject token’s logit at the middle layers. To a lesser degree, it also promotes answer tokens, particularly at layer 2525.888Thus, the observation from Section 4.2 that attention promotes answers at layer $25$ can be attributed to the subject token. Answer promotion is most pronounced in the Country-Cities dataset (Figure˜24, Figure˜25) but less evident in the Artist-Songs and Actor-Movies datasets (Figure˜26, Figure˜27, Figure˜28, Figure˜29). In all datasets, the subject logit is still larger than that of each answer across all answer steps, demonstrating that attention primarily copies or propagates subject information from the subject to the last token position.
@@ 6.3 Role of Previous Answer Tokens
As shown in Figure˜5, all previous answer tokens have negative MLP logit differences at late layers. For instance, $o_{1}$ has negative logits at step $2$ starting around layer $27$; $o_{1}$ and $o_{2}$ exhibit similar patterns at step $3$. This suppression aligns with attention’s role in inhibiting previously generated tokens, suggesting that MLPs amplify these suppression signals to prevent repetition.
@@
Surprisingly, we observe positive MLP logit differences for new answers across answer steps (Figure˜5). Specifically, the logit differences of both $o_{2}$ and $o_{3}$ are positive when intervening on $o_{1}$ at step $2$; $o_{3}$ has positive logits differences when intervening on $o_{1}$ or $o_{2}$ at step $3$. This pattern shows that MLPs also leverage previous answer tokens to promote new answers. Since LMs already promote all relevant answers when predicting previous answers, it is plausible that the models reuse these prior computations to promote new answers. These findings show that the subject token is not the sole source of answer promotion (Section˜5.1). The previous answer tokens also have a positive (but smaller) effect on answer promotion.
@@ 6.4 Role of Last Token
Starting from layer $28$, attention from the last token to itself significantly increases the logit of all three answers (Figure˜6). At each step $i$, the logit for the answer $o^{(i)}$} is consistently the highest among the three answers. This result suggests that attention at the last token aggregates information from earlier layers related to knowledge recall and suppression, preparing the model for generating the next prediction.
@@
Interestingly, we observe MLPs output negative logit differences for the subject and all three answers in the final layers when knocking out the attention from the last token to itself (Figure˜6). The answer $o^{(i)}$} for each step $i$ consistently has the most negative logit differences. In other words, without having access to the attention output of the last token, MLPs output even higher logits for the subject and the answers. This behavior suggests a backup mechanism: without direct attention to the last token that aggregates information from early input tokens, the model may not have sufficient promotion and differentiation of the three answers. MLPs compensate this by further promoting the three answers to encourage the predictions to be correct. Similarly, Wang et al. (2022) find backup token mover attention heads that become active when the original token mover heads are ablated.
7 Are Knowledge Recall and Suppression Independent?